# An ultraslow optical centrifuge with arbitrarily low rotational acceleration

### v1

@@ Abstract
We outline the design and characterization of a laser pulse shaper, which creates an “ultraslow optical centrifuge” - a linearly polarized field whose polarization vector rotates with arbitrarily low angular acceleration. By directly recording this rotation in time with nonlinear cross-correlation, we demonstrate the tunability of such centrifuge (both in terms of its initial and its final rotational frequencies) in the range of accelerations which are three orders of magnitude lower than those available with a conventional centrifuge design. We showcase the functionality of the ultraslow centrifuge by spinning css molecules in a molecular jet. Utilizing the extremely low angular acceleration to control molecular rotation inside viscous media is a promising application for this unique optical tool.
@@ I Introduction
In a recent publication [1], we argued that the ability of an optical centrifuge – an ultrafast laser-based instrument with numerous applications in molecular sciences (for recent reviews, see Refs. 2, 3) – to control molecular rotation is limited to molecules with relatively low moment of inertia and centrifugal distortion constant. The former sets an upper bound to the angular acceleration of the centrifuge, whereas the latter constrains the range of accessible rotational frequencies. By the very design of the original optical centrifuge [4, 5], it provided an efficient way to spin up molecules in the gas phase at a fast rate of $\sim$100\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}\mathrm{/}\mathrm{p}\mathrm{s}$\$ to extremely high rotational states, known as “super-rotors”, reaching frequencies of order $10\text{\,}\mathrm{T}\mathrm{H}\mathrm{z}$.
@@
In Ref. [1], we reported on the novel design of an optical centrifuge, based on the concept of spectral focusing [17], which produced zero angular acceleration. This “constant-frequency” centrifuge (cfCFG) allowed us to operate in the desired low range of rotational frequencies around $10\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$, and has already proven capable of controlled rotational excitation of molecules embedded in helium nanodroplets [18]. At the same time, zero acceleration of cfCFG also meant that the enforced adiabatic rotational ladder climbing (typically expected from a conventional optical centrifuge) could not be executed, which resulted in a rather limited degree of rotational control.
@@ II Concept of ultraslow acceleration
After delaying one centrifuge arm with respect to the other by $\Delta t$, their interference results in the constant angular frequency of cfCFG,
@@
where we neglected the term $\Delta\beta\Delta t/2$, since $|\Delta\beta|\ll|\beta_{0}|$. Eq. (4) above describes an ultraslow optical centrifuge with constant acceleration
$\varepsilon_{\mathrm{us}}=\frac{\dot{\Omega}_{\mathrm{us}}(t)}{2\pi}=\frac{\Delta\beta}{2\pi}.$ (5)
@@
The time-dependent centrifuge frequency can be expressed as:
$f_{\mathrm{us}}(t)=\frac{\Omega_{\mathrm{us}}(t)}{2\pi}=f_{0}+\varepsilon_{\mathrm{us}}t,$ (6)
@@
$\Delta f_{\mathrm{us}}=\varepsilon_{\mathrm{us}}\tau=\frac{\Delta\beta\tau}{2\pi},$ (8)
@@
Figure 2: Ultraslow centrifuge pulse shaper. The input pulse (purple) is split into two arms with a half-wave plate ($\lambda/2$) and a polarizing beam splitter (PBS). The transmitted portion (green) is sent through the double-grating pulse compressor with a variable distance $L$ between the gratings ($\mathrm{GR}_{1}$ and $\mathrm{GR}_{2}$). After being vertically displaced with a retro-reflector ($R$), this “compressor arm” is overlapped with the “delay arm” (cyan) on the output PBS. IM/OM mark the input/output mirrors of the compressor, respectively. A quarter-wave plate ($\lambda/4$) converts the two orthogonal linear polarizations into the circular polarizations of opposite handedness, to form the field of the centrifuge (corkscrew shape). The inset shows our definition of the grating orientation angle $\theta_{0}$, used in the main text.
@@ III Ultraslow centrifuge shaper
As a relevant example of usCFG parameters, consider the centrifuge, which would accelerate from 0 to $30\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$. This implies a central frequency $f_{0}=$15\text{\,}\mathrm{GHz}$\$ and a bandwidth $\Delta f_{\mathrm{us}}=$30\text{\,}\mathrm{GHz}$\$. From the parameters of our laser pulses, their central wavelength $\lambda_{0}=$803\text{\,}\mathrm{nm}$\$, spectral width $\Delta\lambda=$8.2\text{\,}\mathrm{n}\mathrm{m}$\$ and pulse length $\tau\approx$320\text{\,}\mathrm{ps}$\$ (both expressed as full width at half maximum), the initial chirp coefficient can be estimated as
@@ IV Calibration and tuning techniques
The cross-correlation is performed by projecting the centrifuge field onto a single polarization axis with a linear polarizer, and then mixing it with a much shorter ($\approx 120\text{\,}\mathrm{fs}$) reference pulse on a nonlinear crystal (barium borate, BBO), as depicted in Fig. 3. To determine the usCFG polarization angle as a function of time, we scan the delay between the reference and the centrifuge pulse, while recording the sum-frequency generation (SFG) signal. As the latter is proportional to the square of the centrifuge field projection, the SFG signal is expected to oscillate at twice $f_{\mathrm{us}}{}(t)$.
Figure 4: (a) Sample cross-correlation trace of usCFG (red diamonds), and its reconstruction from the numerically extracted instantaneous frequency (solid black). The dashed black line outlines the pulse envelope. (b) Extracted instantaneous frequency (red), where positive and negative values represent opposite directions of rotation. The dashed black curve shows $f_{\mathrm{us}}(t)$, modeled to minimize the reconstruction error of the cross-correlation trace from the solid black line (a), with quadratic time dependence to account for third-order dispersion (see text for details).
@@
The rotation decelerates in the first half of the pulse, reaching zero frequency at time zero, and then accelerates in the opposite direction. We extract the absolute value of the instantaneous frequencies of rotation by calculating the Hilbert transform of the cross-correlation signal, and differentiating the phase of the complex-valued result with respect to time. Negative sign is applied to the frequencies on the left of the inflection point, if the latter is detected. After smoothing the extracted frequency-vs-time dependence, we use it as an initial guess for fitting the signal with a frequency-chirped squared sinusoid oscillating under a Gaussian envelope. Time $t=0$ is assigned to the peak of that envelope. The corresponding rotational frequency of the centrifuge (equal to half the oscillation frequency of the SFG signal) is plotted in Fig. 4(b), showing the desired (almost) constant acceleration across the centrifuge profile.
A slight increase in the rotational acceleration with time, apparent in Fig. 4(b), can be attributed to third-order dispersion (TOD), present in the initial chirped pulses from the laser system ($\gamma_{0}$; the effect of this TOD on the centrifuge acceleration has been analyzed in detail in Ref. 1), as well as added by the compressor arm of the centrifuge shaper ($\Delta\gamma$). Owing to such TOD, the optical phase is varying in time as
@@
Figure 5: Tunable behaviour of usCFG, characterized by means of optical cross-correlation. (a) Central frequency of usCFG, $f_{0}$, at different delays between the centrifuge arms, $\Delta t$. The grating separation in the compressor was recorded with a fixed $L\approx$39\text{\,}\mathrm{mm}$\$. (b) The frequency bandwidth of usCFG, $\Delta f_{\mathrm{us}}$, at different grating separations $L$, recorded with a fixed value of $\Delta t\approx$3\text{\,}\mathrm{ps}$\$. (c) Measurements of $f_{\mathrm{us}}(t)$ for three different centrifuges, tuned to have terminal frequencies of approximately $20\text{\,}\mathrm{GHz}$ (solid blue), $30\text{\,}\mathrm{GHz}$ (dashed orange), and $40\text{\,}\mathrm{GHz}$ (dotted green). By adjusting both $\Delta t$ and $L$ (and, thus, $f_{0}$ and $\Delta f_{\mathrm{us}}{}$) in a coordinated fashion, all centrifuges were tuned to pass zero frequency around $150\text{\,}\mathrm{ps}$, but reach different terminal frequencies by the end of the pulse at $150\text{\,}\mathrm{ps}$.
To illustrate the tunability of the central frequency of the ultraslow centrifuge, we performed cross-correlation measurements for various values of $\Delta t$, while keeping $\Delta\beta$ constant. Fig. 5(a) shows the expected linear relationship between $f_{0}$ and $\Delta t$, as predicted by Eq. 7. The slope of $f_{0}/\Delta t=\beta_{0}/2\pi=$6.30(2)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{ps}}^{-1}$\$ is similar to the value of $6.4(2)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{ps}}^{-1}$ reported in Ref [1] for a constant-frequency centrifuge constructed from the same laser source (and, therefore, with the same $\beta_{0}$).
To demonstrate our ability to vary the acceleration of usCFG, Fig. 5(b) shows the extracted values of $\Delta f_{\mathrm{us}}{}$ at different values of the grating separation $L$, recorded with fixed $\Delta t\approx$3\text{\,}\mathrm{ps}$\$. Here again, we observe a linear dependence, as anticipated from Equations (8) and (9). The slope of $0.70(2)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{mm}}^{-1}$ is in good agreement with the calculated value of $0.71(1)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{mm}}^{-1}$.
Despite the effects of TOD on the exact characteristics of usCFG, and the precise time-dependence of $f_{\mathrm{us}}(t)$, the key centrifuge parameters can still be easily controlled. In Fig. 5(c), we illustrate this by adjusting the terminal frequency of usCFG to three different values, while keeping the initial frequency close to zero, as is often required for the adiabatic spinning of cold molecules.
@@ V Proof of operation in the molecular jet
To demonstrate the utility of usCFG to induce molecular rotation, we employed the well-known technique of Velocity Map Imaging (VMI) [20]. Centrifuge pulses are focused onto a molecular jet of carbon disulfide (css) molecules, which are then Coulomb-exploded by a $\approx 120\text{\,}\mathrm{fs}$ probe pulse. css gas is mixed with $8\text{\,}\mathrm{bar}$ of helium at a concentration of $1000\text{\,}\mathrm{p}\mathrm{p}\mathrm{m}$, with an estimated rotational temperature below $10\text{\,}\mathrm{K}$ after supersonic expansion through a pulsed nozzle into vacuum.
@@
Figure 6: Geometry of the Velocity Map Imaging experiment. The magenta corkscrews represent the electric field of the centrifuge ($\vec{E}_{\mathrm{CFG}}$), propagating along the $Y$ axis. The VMI setup is oriented along the $Z$ axis, projecting a two-dimensional velocity map image of ion fragments onto the detector, shown below the XYXY plane. $\theta_{\text{2D}}$ is the angle of a detected ion fragment, measured with respect to the XZXZ plane. A css molecule is shown aligned to $\vec{E}_{\mathrm{CFG}}$.
Fig. 7(a) shows the degree of alignment of css molecules during the usCFG. The centrifuge was tuned to start at zero rotational frequency, i.e. $f_{\mathrm{us}}{}(t\approx$-150\text{\,}\mathrm{p}\mathrm{s}$)=$0\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$\$, and accelerate to $\approx 25\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$ at the peak of its intensity profile around $t=0$ ps. Clear oscillations of $\langle\cos^{2}\theta_{\text{2D}}\rangle{}(t)$ confirm that css molecules have been captured by the centrifuge and follow its rotation. As molecules spin in the $XZ$ plane, the ejection angle of S+ ions periodically alternates between along the $X$-axis, resulting in $\langle\cos^{2}\theta_{\text{2D}}\rangle>0.5$, and along the $Z$-axis, with $\langle\cos^{2}\theta_{\text{2D}}\rangle\approx 0.5$. Evidently, the oscillation period is decreasing with time, attesting to the rotational acceleration of molecules by usCFG. The spectrogram in Fig. 7(b) was produced by applying a short-time Fourier transform to the data presented in Fig. 7(a), allowing for a time-frequency map of molecular rotation.
Figure 7: (a) $\langle\cos^{2}\theta_{\text{2D}}\rangle(t)$ of S+ fragments (blue markers with red error bars), recorded during the forced rotation of css molecules by usCFG. Blue lines connect the data points to guide the eye. (b) spectrogram produced by applying a short-time Fourier transform to the signal plotted in (a). The tilted trace accelerating up to $\sim$100\text{\,}\mathrm{GHz}$\$ confirms that the molecules are following the accelerated rotation of the centrifuge.

### v2

@@ Abstract
We present the design and characterization of a laser pulse shaper, which creates an “ultraslow optical centrifuge” - a linearly polarized field whose polarization vector rotates with arbitrarily low angular acceleration. By directly recording this rotation in time with nonlinear cross-correlation, we demonstrate the tunability of the initial and final rotational frequencies of such a centrifuge in the range of accelerations, which are three orders of magnitude lower than those available with a conventional centrifuge design. We showcase the functionality of the ultraslow centrifuge by spinning css molecules in a molecular jet. Utilizing the extremely low angular acceleration to control molecular rotation inside viscous media is a promising application for this unique optical tool.
@@ I Introduction
An optical centrifuge is an ultrafast laser-based instrument with numerous applications in molecular sciences (for recent reviews, see Refs. 1, 2). In a recent publication [3], we argued that the ability of such a device to control molecular rotation is limited to molecules with relatively low moment of inertia and centrifugal distortion constant. The former sets an upper bound to the angular acceleration of the centrifuge, whereas the latter constrains the range of accessible rotational frequencies. By the very design of the original optical centrifuge [4, 5], it provided an efficient way to spin up molecules in the gas phase at a fast rate of $\sim$100\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}\mathrm{/}\mathrm{p}\mathrm{s}$\$ to extremely high rotational states, known as “super-rotors”, reaching frequencies of order $10\text{\,}\mathrm{T}\mathrm{H}\mathrm{z}$.
@@
An alternative approach to creating an optical centrifuge with zero angular acceleration by using two time-delayed frequency-chirped laser pulses of opposite circular polarization has been proposed theoretically [17], and used as a method of frequency shearing in a transparent dielectric medium for ultrashort pulse characterization [18]. Applying this method to molecular ensembles, in Ref. [3] we demonstrated experimentally that such a “constant-frequency” centrifuge (cfCFG) is indeed capable of controlling molecular rotation in the desired low range of rotational frequencies around $10\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$, finally enabling us to optically centrifuge molecules embedded in helium nanodroplets [19]. At the same time, zero acceleration of cfCFG also meant that the enforced adiabatic rotational ladder climbing (typically expected from a conventional optical centrifuge) could not be executed, which resulted in a rather limited degree of rotational control.
@@ II Concept of ultraslow acceleration
The angular frequency of the polarization vector is half the difference in optical frequencies between the two arms. After delaying one centrifuge arm with respect to the other by $\Delta t$, their interference results in the constant angular frequency of cfCFG,
@@
where we neglected the term $\Delta\beta\Delta t/2$, since $|\Delta\beta|\ll|\beta_{0}|$. Eq. (4) above describes an ultraslow optical centrifuge with constant angular acceleration $\dot{\Omega}_{\mathrm{us}}=\Delta\beta$.
@@
The time-dependent rotational frequency of the centrifuge can be expressed as:
$f_{\mathrm{us}}(t)=\frac{\Omega_{\mathrm{us}}(t)}{2\pi}=f_{0}+\frac{\Delta\beta}{2\pi}t,$ (5)
@@
$\Delta f_{\mathrm{us}}=\frac{\Delta\beta\tau}{2\pi},$ (7)
@@
Figure 2: Ultraslow centrifuge pulse shaper. The chirped input pulse (purple) is split into two arms with a half-wave plate ($\lambda/2$) and a polarizing beam splitter (PBS). The transmitted “compressor arm” (green beam) is sent through the double-grating pulse compressor with a variable distance $L$ between the gratings ($\mathrm{GR}_{1}$ and $\mathrm{GR}_{2}$), controlled with a translation stage (green rectangle), and is vertically displaced with a retro-reflector ($R$). Another translation stage (blue rectangle) controls the delay $\Delta t$ of the reflected “delay arm” (cyan beam). The two arms are combined on the output PBS. IM/OM mark the input/output mirrors of the compressor, respectively. A quarter-wave plate ($\lambda/4$) converts the two orthogonal linear polarizations into the circular polarizations of opposite handedness, to form the field of the centrifuge (corkscrew shape). The inset shows our definition of the grating orientation angle $\theta_{0}$, used in the main text.
@@ III Ultraslow centrifuge shaper
As a relevant example of usCFG parameters, consider the centrifuge, which would accelerate from 0 to $30\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$. This implies a central frequency $f_{0}=$15\text{\,}\mathrm{GHz}$\$ and a bandwidth $\Delta f_{\mathrm{us}}=$30\text{\,}\mathrm{GHz}$\$. From the parameters of our laser pulses, their central wavelength $\lambda_{0}=$802\text{\,}\mathrm{nm}$\$, spectral width $\Delta\lambda=$8.2\text{\,}\mathrm{n}\mathrm{m}$\$ and pulse length $\tau\approx$320\text{\,}\mathrm{ps}$\$ (both expressed as full width at half maximum), the initial chirp coefficient can be estimated as
@@ IV Calibration and tuning techniques
The cross-correlation is performed by projecting the centrifuge field onto a single polarization axis with a linear polarizer, and then mixing it with a much shorter ($\approx 120\text{\,}\mathrm{fs}$) reference pulse of the same polarization on a type-I nonlinear crystal (barium borate, BBO), as depicted in Fig. 3. To determine the usCFG polarization angle as a function of time, we scan the delay between the reference and the centrifuge pulse, while recording the sum-frequency generation (SFG) signal. As the latter is proportional to the square of the centrifuge field projection, the SFG signal is expected to oscillate at twice $f_{\mathrm{us}}{}(t)$.
Figure 4: (a) Sample cross-correlation trace of usCFG (red diamonds), and its reconstruction from the numerically extracted instantaneous frequency (solid black). The dashed black line outlines the pulse envelope. (b) Extracted instantaneous frequency smoothed with a moving average (red). The dashed black curve shows $f_{\mathrm{us}}(t)$, modeled to minimize the reconstruction error of the cross-correlation trace from the solid black line (a), with quadratic time dependence to account for third-order dispersion (see text for details).
@@
The rotation decelerates in the first half of the pulse, reaching zero frequency at time zero, and then accelerates in the opposite direction. We extract the absolute value of the instantaneous frequencies of rotation by calculating the Hilbert transform of the cross-correlation signal, and differentiating the phase of the complex-valued result with respect to time. Negative sign is applied to the frequencies on the left of the inflection point, if the latter is detected. After smoothing the extracted frequency-vs-time dependence, we use it as an initial guess for fitting the signal with a frequency-chirped squared sinusoid oscillating under a Gaussian envelope. Time $t=0$ is assigned to the peak of that envelope. The corresponding rotational frequency of the centrifuge (equal to half the oscillation frequency of the SFG signal) is plotted in Fig. 4(b), showing near-constant acceleration across the centrifuge profile.
A slight increase in the rotational acceleration with time is apparent in Fig. 4(b) and can be attributed to positive third-order dispersion (TOD) in the initial chirped pulses from the laser system ($\gamma_{0}$; the effect of this TOD on the centrifuge acceleration has been analyzed in detail in Ref. 3) and that added by the grating pair, $\Delta\gamma$. Owing to such TOD, the optical phase is varying in time as
@@
Figure 5: Tunable behaviour of usCFG, characterized by means of optical cross-correlation. (a) Central frequency of usCFG, $f_{0}$, at different delays between the centrifuge arms, $\Delta t$. The grating separation in the compressor was recorded with a fixed $L\approx$39\text{\,}\mathrm{mm}$\$. (b) The frequency bandwidth of usCFG, $\Delta f_{\mathrm{us}}$, at different grating separations $L$, recorded with a fixed value of $\Delta t\approx$3\text{\,}\mathrm{ps}$\$. (c) Extracted $f_{\mathrm{us}}(t)$ smoothed with a moving average for three different centrifuges, tuned to have terminal frequencies of approximately $20\text{\,}\mathrm{GHz}$ (blue), $30\text{\,}\mathrm{GHz}$ (orange), and $40\text{\,}\mathrm{GHz}$ (green).
To illustrate the tunability of the central frequency of the ultraslow centrifuge, we performed cross-correlation measurements for various values of $\Delta t$, while keeping $L$ constant. Fig. 5(a) shows the expected linear relationship between $f_{0}$ and $\Delta t$, as predicted by Eq. (6), and the desired minimal changes in $\Delta f_{\mathrm{us}}{}$. The slope $f_{0}/\Delta t=\beta_{0}/2\pi=$6.30(2)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{ps}}^{-1}$\$ is similar to the value of $6.4(2)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{ps}}^{-1}$ reported in Ref [3] for a cfCFG constructed from the same laser source with the same $\beta_{0}$.
To demonstrate our ability to vary the acceleration of usCFG, Fig. 5(b) shows the extracted values of $\Delta f_{\mathrm{us}}{}$ at different values of the grating separation $L$, recorded with fixed delay stage position. Here again, we observe a linear dependence, as anticipated from Equations (7) and (8). The slope of $0.70(2)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{mm}}^{-1}$ is in good agreement with the calculated value of $0.71(1)\text{\,}\mathrm{GHz}\text{\,}{\mathrm{mm}}^{-1}$. Owing to the imperfections of the beam alignment in the centrifuge shaper, slight unintentional changes in the delay between the two interferometer arms always accompany the deliberate tuning of the grating separation, resulting in small variations of $f_{0}$.
Despite the nonlinear effects of TOD on $f_{\mathrm{us}}(t)$ and the calibration characteristics of usCFG, tuning the centrifuge is readily accessible by appropriately adjusting both $\Delta t$ and $L$ (and thus $f_{0}$ and $\Delta f_{\mathrm{us}}{}$). In Fig. 5(c), we demonstrate the performances of three different centrifuges that all start from around zero frequency at the rising edge of the pulse, but reach different terminal frequencies by its falling edge.
@@ V Proof of operation in the molecular jet
[…]
@@
Figure 6: Geometry of the Velocity Map Imaging experiment. The magenta corkscrews represent the electric field of the centrifuge ($\vec{\mathbf{E}}_{\mathrm{CFG}}$), propagating along the $Y$ axis. The VMI setup is oriented along the $Z$ axis, projecting a two-dimensional velocity map image of ion fragments onto the detector, shown below the XYXY plane. $\theta_{\text{2D}}$ is the angle of a detected ion fragment, measured with respect to the XZXZ plane. A css molecule is shown aligned to $\vec{\mathbf{E}}_{\mathrm{CFG}}$.
Fig. 7(a) shows the degree of alignment of css molecules during the usCFG. The centrifuge was tuned to start at zero rotational frequency, i.e. $f_{\mathrm{us}}{}(t\approx$-150\text{\,}\mathrm{p}\mathrm{s}$)=$0\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$\$, and accelerate to $\approx 25\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$ at the peak of its intensity profile around $t=0$ ps. Given the initial chirp rate $\beta_{0}\approx$ $40\text{\,}\mathrm{rad}\,\mathrm{p}\mathrm{s}^{-2}$, this required the following parameters of the centrifuge shaper: $\Delta\beta\approx$ $1\text{\times}{10}^{-3}\text{\,}\mathrm{rad}\,\mathrm{p}\mathrm{s}^{-2}$, and $\Delta t\approx$ $4\text{\,}\mathrm{ps}$.
Clear oscillations of $\langle\cos^{2}\theta_{\text{2D}}\rangle{}(t)$ confirm that css molecules have been captured by the centrifuge and follow its rotation. As molecules spin in the $XZ$ plane, the ejection angle of S+ ions periodically alternates between along the $X$-axis, resulting in $\langle\cos^{2}\theta_{\text{2D}}\rangle>0.5$, and along the $Z$-axis, with $\langle\cos^{2}\theta_{\text{2D}}\rangle\approx 0.5$. Evidently, the oscillation period is decreasing with time, attesting to the rotational acceleration of molecules by usCFG. The spectrogram in Fig. 7(b) was produced by applying a short-time Fourier transform to the data presented in Fig. 7(a), allowing for a time-frequency map of molecular rotation.
Figure 7: (a) $\langle\cos^{2}\theta_{\text{2D}}\rangle(t)$ of S+ fragments (blue markers with red error bars), recorded during the forced rotation of css molecules by usCFG. Blue lines connect the data points to guide the eye. (b) Spectrogram produced by applying a short-time Fourier transform (STFT) to the signal plotted in (a). The tilted trace accelerating up to $\sim$100\text{\,}\mathrm{GHz}$\$ confirms that the molecules are following the accelerated rotation of the centrifuge. The equidistant lines are numerical artifacts related to the window size of the STFT procedure.